Hypatia could row 90 miles downstream in 6 hours but required 8 hours to row 72 miles upstream. What was her speed in still water and what was the speed of the current?
Answer by checkley77(12844)

D=RT
90=R*6
R=90/6
R=15 MPH.
72=R*8
R=72/8
R=9 MPH.
(15+9)/2=24/2=12 MPH. IS THE RATE OF BOAT IN STILL WATER.
(15-9)/2=6/2=3 MPH. IS THE RATE OF THE STREAM.
PROOF:
90=(12+3)*6
90=15*6
90=90
72=(12-3)*8
72=9*8
72=72